we wist to select two  person ffom 8 person but if A is choosen then B must be choosen in how many ways can the selection be made
 

Grade:11

1 Answers

Arun
6 years ago
 

case 1  :   A and B is not included

 

 then number of ways = 6C2 = 15

 

 

 

case 2 : B is selected A is not selected

 

 number of ways  = 1* 6C1  = 6

 

 

 

case 3:  A and B both are selected

 

 number of ways = 1*1 = 1

 

 

 

so total ways  = 1+6+15 =22